# Contextual self-paced learning for Weakly Supervised Spatio-Temporal Video Grounding

### v1

@@ 1 Introduction
Weakly supervised learning has been studied extensively in the image domain, addressing tasks like phrase grounding Datta et al. (2019); Wang et al. (2020); Liu et al. (2021) and referral grounding Liu et al. (2019; 2022b), which locate objects in images based on text. Various methods have been explored, such as those leveraging coarse image-level labels or proposing complex mechanisms to handle uncertainty in object localization. However, extending these approaches to videos adds a new layer of complexity due to dynamic changes in subject poses and scene context over time. As shown in Figure 1, STVG involves increased complexity compared to static image tasks, particularly when handling free-form textual queries, where models must understand and localize objects and actions described in natural language. Existing works that address WSTVG rely on computationally expensive solutions, such as hierarchical algorithms Li et al. (2023) or the inclusion of extra modality data like optical flow Chen et al. (2019b). In contrast, we propose a more streamlined and efficient approach that simplifies the process by focusing solely on visual and textual modalities.
@@
We build upon recent progress in multimodal learning and leverage vision-language foundation models as our baseline, specifically adapting Grounding DINO (G-DINO) Liu et al. (2023), a model known for its strong zero-shot capabilities in image-level tasks. While this model shows promise for multimodal understanding, extending it to STVG reveals three key limitations (Table 1). First, it struggles with temporal consistency, frequently switching object focus across frames, as it lacks a clear understanding of temporal grounding. Second, despite being trained on large-scale image-text datasets, it finds it difficult to handle complex or imbalanced queries, particularly when multiple objects or activities are described simultaneously. Finally, the model’s performance declines in dense scenes with numerous objects, where accurate localization becomes critical.
@@ 2 Related Work
Spatio-Temporal Video Grounding: This task involves grounding spatio-temporal tubes based on textual queries, addressing spatial and temporal dimensions. Initial solutions use a two-stage process with separate spatial (Rohrbach et al., 2015; Yamaguchi et al., 2017; Chen et al., 2019c) and temporal grounding (Gao et al., 2017; Chen et al., 2019a). However, pre-trained object detectors have a fixed number of object categories, limiting their effectiveness for STVG tasks with free-form queries. Recent multimodal approaches (Su et al., 2021; Yang et al., 2022; Jin et al., 2022; Lin et al., 2023; Gu et al., 2024; Wasim et al., 2024) tackle this challenge in a single stage, leveraging image-based detectors (Kamath et al., 2021), video encoders, and spatio-temporal decoders (Yang et al., 2022; Wasim et al., 2024), addressing feature alignment inconsistencies (Jin et al., 2022), or utilizing static and motion cues (Lin et al., 2023; Gu et al., 2024), . These methods typically rely on frame-level bounding box annotations for training. Differently from these, our work adopts a cost-efficient approach by refraining from using spatio-temporal labels.
Weakly Supervised Learning For grounding techniques, it can be categorized into three main classes. In images, it includes phrase and referral grounding. Phrase grounding(Rohrbach et al., 2016; Datta et al., 2019; Chen et al., 2018; Akbari et al., 2019; Gupta et al., 2020; Wang et al., 2020; Liu et al., 2021; Wang et al., 2021a) highlights objects in textual queries using margin losses (Datta et al., 2019; Chen et al., 2018), contrastive optimization (Gupta et al., 2020; Wang et al., 2020), and reconstruction (Rohrbach et al., 2016) methods. Referral grounding (Liu et al., 2019; 2022b; Jin et al., 2023) adopts reconstruction (Liu et al., 2019; 2022b) or contrastive learning (Jin et al., 2023) to ground objects. In temporal grounding for videos (Wang et al., 2021b; Chen et al., 2022; Lin et al., 2020; Zheng et al., 2022a; b), both reconstruction and contrastive methods are prominent, however recent reconstruction-based approaches (Lin et al., 2020; Zheng et al., 2022a; b) outperform contrastive ones. We employ a contrastive and reconstructive approach for spatial and temporal grounding respectively. Different from existing works, we incorporate referential capabilities in spatial and temporal grounding for videos which previous work don’t. Our approach induce focusing on specific contextual knowledge to enhance mutual interaction between vision and text.
@@ 5.1 Ablation Study
Table 6 shows CoSPaL outperforms WINNER with Faster R-CNN Anderson et al. (2018) backbone for fair comparison. Comparing across backbones, DETR outperforms Faster R-CNN by a margin of 6% at m_vIoU on HCSTVG-v1.
@@
Impact on actor localization: Compared to W-GDINO and TPG, CoSPaL improve the classification accuracy by a margin of 20% and 3.2% respectively on HCSTVG-v1 dataset.

### v3

@@ 1 Introduction
Weakly supervised learning has been studied extensively in the image domain, addressing tasks like phrase grounding (Datta et al., 2019; Wang et al., 2020; Liu et al., 2021) and referral grounding (Liu et al., 2019; 2022b), which locate objects in images based on text. Various methods have been explored, such as those leveraging coarse image-level labels or proposing complex mechanisms to handle uncertainty in object localization. However, extending these approaches to videos adds a new layer of complexity due to dynamic changes in subject poses and scene context over time. As shown in Figure 1, STVG involves increased complexity compared to static image tasks, particularly when handling free-form textual queries, where models must understand and localize objects and actions described in natural language. Existing works that address WSTVG rely on computationally expensive solutions, such as hierarchical algorithms (Li et al., 2023) or the inclusion of extra modality data like optical flow (Chen et al., 2019b). In contrast, we propose a more streamlined and efficient approach that simplifies the process by focusing solely on visual and textual modalities.
@@
We build upon recent progress in multimodal learning (Madan et al., 2024) and leverage vision-language foundation models as our baseline, specifically adapting Grounding DINO (G-DINO) (Liu et al., 2023), a model known for its strong zero-shot capabilities in image-level tasks. While this model shows promise for multimodal understanding, extending it to STVG reveals three key limitations (Table 1). First, it struggles with temporal consistency, frequently switching object focus across frames, as it lacks a clear understanding of temporal grounding. Second, despite being trained on large-scale image-text datasets, it finds it difficult to handle complex or imbalanced queries, particularly when multiple objects or activities are described simultaneously. Finally, the model’s performance declines in dense scenes with numerous objects, where accurate localization becomes critical.
@@ 2 Related Work
Spatio-Temporal Video Grounding: This task involves grounding spatio-temporal tubes based on textual queries, addressing spatial and temporal dimensions. Initial solutions developed for STVG use a two-stage process with separate spatial (Rohrbach et al., 2015; Yamaguchi et al., 2017; Chen et al., 2019c) and temporal grounding (Gao et al., 2017; Chen et al., 2019a). However, pre-trained object detectors have a fixed number of object categories, limiting their effectiveness for STVG tasks with free-form queries. Recent multimodal approaches (Su et al., 2021; Yang et al., 2022; Jin et al., 2022; Lin et al., 2023; Gu et al., 2024; Wasim et al., 2024) tackle this challenge in a single stage, leveraging image-based detectors (Kamath et al., 2021), video encoders, and spatio-temporal decoders (Yang et al., 2022; Wasim et al., 2024), addressing feature alignment inconsistencies (Jin et al., 2022), or utilizing static and motion cues (Lin et al., 2023; Gu et al., 2024), . These methods typically rely on frame-level bounding box annotations for training. Differently from these, our work adopts a cost-efficient approach by refraining from using spatio-temporal labels.
Weakly Supervised Learning There are some existing works on dense tasks (Kumar et al., 2024; Singh et al., 2024; Kumar et al., 2023; Rana & Rawat, 2023; Kumar & Rawat, 2022; Rana & Rawat, 2022; Dave et al., 2022; Modi et al., 2022), however, they are unimodal and on semi-supervised or active learning and can’t be extended to solve weakly supervised STVG task. For grounding techniques, it can be categorized into three main classes. In images, it includes phrase and referral grounding. Phrase grounding(Rohrbach et al., 2016; Datta et al., 2019; Chen et al., 2018; Akbari et al., 2019; Gupta et al., 2020; Wang et al., 2020; Liu et al., 2021; Wang et al., 2021a) highlights objects in textual queries using margin losses (Datta et al., 2019; Chen et al., 2018), contrastive optimization (Gupta et al., 2020; Wang et al., 2020), and reconstruction (Rohrbach et al., 2016) methods. Referral grounding (Liu et al., 2019; 2022b; Jin et al., 2023) adopts reconstruction (Liu et al., 2019; 2022b) or contrastive learning (Jin et al., 2023) to ground objects. In temporal grounding for videos (Wang et al., 2021b; Chen et al., 2022; Lin et al., 2020; Zheng et al., 2022a; b), both reconstruction and contrastive methods are prominent, however recent reconstruction-based approaches (Lin et al., 2020; Zheng et al., 2022a; b) outperform contrastive ones. We employ a contrastive and reconstructive approach for spatial and temporal grounding respectively. Different from existing works, we incorporate referential capabilities in spatial and temporal grounding for videos which previous work don’t. Our approach induce focusing on specific contextual knowledge to enhance mutual interaction between vision and text.
@@ 5.1 Ablation Study
Table 6 shows CoSPaL outperforms WINNER with Faster R-CNN (Anderson et al., 2018) backbone for fair comparison. Comparing across backbones, DETR outperforms Faster R-CNN by a margin of 6% at m_vIoU on HCSTVG-v1.
@@ 6 Conclusion
7 Acknowledgments
This research is based upon work supported in part by the Office of the Director of National Intelligence (Intelligence Advanced Research Projects Activity) via 2022-21102100001 and the National Science Foundation under Grant No. 2239292. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the US Government. The US Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein. The authors would like to thank Aisha Urooj Khan (Lunit) and Rohit Gupta (UCF CRCV) for helpful discussions on weakly supervised grounding and multimodal aspects.